# MACE-OFF23: Transferable Machine Learning Force Fields for Organic Molecules

### v2

@@ III.1 Extended SPICE Test Set
First, we look at the pointwise errors of the energy and force predictions on a held-out test set for each of the three MACE-OFF models. Figure 1 shows the per-atom energy, force and intermolecular force root-mean-square errors (full statistics are provided in SI Table 3). As the size of the model increases, the models gradually become more accurate, with the large model generally achieving errors of around 0.5–1.0 meV/atom and 15–20 meV/Å, well below the 1 kcal/mol (43 meV) chemical accuracy limit for the drug-like organic molecules studied here.
Figure 1: Test set root mean square errors (RMSE). Errors in the MACE-OFF23 models compared to the underlying DFT reference data, highlighting the relative accuracy of the three models. Bottom panels show specifically inter-molecular force errors compared to overall DFT inter-molecular force magnitudes (RMS).
The last column looks at an extrapolation task, the training set contains only dipeptides, and this test set looks at tripeptides, indicating that the models are able to extrapolate to larger fragments with more complex interactions, with no degradation of accuracy.
The bottom panel of Figure 1 shows specifically the intermolecular force errors, which were obtained by separating the force contributions to molecular translations and rotations (see SI and Ref. [63] for details). These interactions are crucially important as they underpin the thermodynamics and transport properties in the organic condensed phase. MACE medium and large models yield similar errors of around 5–15 meV/Å, which are about 1.5–3 times smaller than total force errors, and about 5–10% of the typical DFT intermolecular force magnitude (RMS). For context, relative intramolecular force errors, which are roughly equivalent to total force errors, are around 1–2%. The latter are routinely used across the literature to benchmark ML models, however, we advocate that intermolecular RMSE errors are a more appropriate metric to assess model accuracy for organic condensed phase applications. The results here demonstrate the remarkable accuracy of MACE models in predicting these subtle but crucial interactions.
@@ III.2 Dihedral scans
Next, we evaluate the performance of the MACE-OFF23 models on dihedral scans of drug-like molecules. This task is routinely carried out using quantum mechanical methods to create reference data for parameterizing classical empirical force field dihedral terms [64]. The task is particularly challenging, as constrained geometry optimisations can be difficult to converge if the potential energy surface is not smooth, as has been observed in particular for the ANI family of potentials [65].
@@ III.2.1 TorsionNet-500
The top panel of Figure 2 summarises the results for the TorsionNet-500 dataset [67]. This dataset contains torsion drives of 500 different molecules, selected to cover a wide range of pharmaceutically relevant chemical space. The original data set was reported at the B3LYP/6-31G(d) level of DFT theory. For consistency with the SPICE dataset, we recreated the torsion profiles using the DFT setting of SPICE, which is a higher level of theory (SI Section C).
The first panel shows an example of a torsion drive, indicating the complex energy profile that the MACE models are able to capture closely, including geometries far from equilibrium. The centre panel shows the mean barrier height error of a number of representative models, comparing the Sage classical empirical force field [21], a semi-empirical quantum mechanical method GFN2-xTB [23], a recent transferable machine learning force field AIMNet-2 [8], and the three MACE-OFF23 models. Again, systematic improvements in accuracy with MACE model size are observed, with medium and large models, in particular, achieving errors of around 0.25 kcal/mol compared to the reference method. The AIMNet-2 model achieves comparable accuracy to the small MACE-OFF23 model. A similar conclusion can be drawn from the comparison of the molecular geometries by looking at the root mean squared deviation of the atomic positions averaged over the full torsion scans, as indicated by the top right panel of Figure 2. It shows that MACE-optimised geometries have a deviation of about 0.025 Å, meaning they are almost indistinguishable from DFT-optimised structures. It is important to note that the different models were trained to different levels of DFT, which might also contribute to the observed differences.
@@ III.2.2 Biaryl fragments
The DFT torsion drives achieve a mean barrier height error of 0.2 kcal/mol with respect to coupled cluster data, and is therefore the best result theoretically possible for MACE-OFF23 models which were trained on the same DFT level of theory. We find that the large model comes close to this, with a mean absolute error of 0.3 kcal/mol. The medium and small MACE models have barrier height errors of 0.4 and 0.5 kcal/mol, respectively. Remarkably, the small MACE model is significantly more accurate than the next-best non-DFT methods, ANI-1ccx and GFN2-xTB, as illustrated in the bottom-center plot of Figure 2. In particular, unlike MACE-OFF23, coupled cluster reference calculations were used to parameterize the ANI-1ccx and GFN2 models. In the bottom right, we also show the cumulative error distributions to verify that the MACE-OFF23 barrier height errors are not only accurate on average but are also robust, having essentially no outliers.
@@ III.3 Molecular crystals
In the following, we demonstrate the ability of the MACE-OFF23 force fields to simulate the vibrational and thermal properties of molecular crystals.
@@ III.3.1 Vibrational spectroscopy of paracetamol
Here, we test the ability of the MACE-OFF23(S) model to predict the Raman spectrum of the “Form II” polymorph of paracetamol. To compare MACE-OFF23(S) directly with experiments, we rigorously incorporate quantum nuclear and non-Condon effects using our ML-aided framework [68]. We incorporate quantum nuclear effects by fitting an effective potential energy surface [69] (using MACE) to calculate quantum nuclear corrections to the MACE-OFF23(S) model within the path-integral coarse-grained simulations (PIGS) method. To incorporate non-Condon effects, we fit a separate equivariant MACE model to the first-principles polarizability of paracetamol polymorphs (data taken from Refs. 70, 71). The remaining steps of our quantum nuclear simulations mimic an entirely classical calculation involving a $NVE$ molecular dynamics simulation, prediction of the isotropic and anisotropic components of the polarizability tensor, and calculation of their time correlation functions. For further details we refer the reader to Ref. 68.
@@ III.3.2 Lattice enthalphies
[…]
@@ III.4 Water structure and dynamics
A key requirement for bio-organic force fields is a good description of water. Several empirical water models have been developed for this purpose, the most common for all-atom biomolecular simulations being TIP3P. Figure 5 shows that both MACE-OFF23(S) and MACE-OFF23(M) correctly predict the radial distribution function (RDF) of water, performing comparably with both TIP3P and the highly accurate MB-pol water model [76]. These interactions are modelled in a purely local way by MACE, having been trained on small water clusters containing up to 50 water molecules. This result further emphasizes the ability of the model to generalize to bulk simulations, even though the training data only contained non-periodic clusters. In contrast, ANI-2X shows a significant over-structuring of the RDF, likely due to the lack of dispersion correction in the underlying quantum mechanical method used to label the training set.
We further calculated the temperature dependence of the density using NPT molecular dynamics simulations, as shown in Supplementary Figure 11. This is a highly sensitive test for the accuracy of intermolecular interactions [63]. Both the small and medium models over-predict the density to a small extent, which likely originates from the underlying DFT functional or the lack of periodic training data.
@@
Figure 5: Water structure and dynamics. Top: IR, isotropic Raman and anisotropic Raman spectra of water at ambient temperature and density using the MACE-OFF23(S) model for the potential energy surface and a single MACE model of water’s dipole and polarizability [68]. The black curve represents the experiment [78]. The dashed blue curve is obtained from standard (classical) MD, while the solid curve incorporates quantum effects using the PIGS method [69]. Bottom: Radial distribution function of MACE, ANI and TIP3P models. MB-pol traces are reproduced from Ref. [76]
Figure 5: Water structure and dynamics. Top: IR, isotropic Raman and anisotropic Raman spectra of water at ambient temperature and density using the MACE-OFF23(S) model for the potential energy surface and a single MACE model of water’s dipole and polarizability [68]. The black curve represents the experiment [78]. The dashed blue curve is obtained from standard (classical) MD, while the solid curve incorporates quantum effects using the PIGS method [69]. Bottom: Radial distribution function of MACE, ANI and TIP3P models. MB-pol traces are reproduced from Ref. [76]
III.5 Condensed phase properties of organic liquids
Small molecule empirical force fields are fitted to reproduce experimental condensed phase properties including densities and heats of vaporization. For machine learning force fields, this task represents a greater challenge, since the potential is typically trained only using QM reference data for isolated molecules and molecular dimers. From these ab initio data, the models infer the long-range interactions necessary to reproduce bulk properties without ever fitting to them.
@@ III.5 Condensed phase properties of organic liquids
[…]
@@
First, we investigate the predictions of molecular liquid densities under ambient conditions using the medium MACE model shown on the top panel of Figure 6. The MACE-OFF23(M) model achieves a MAE of 0.09 g/cm3, indicating that it can generalize from the intermolecular interactions between molecular dimers in the training set to the condensed phase, retaining a good correlation with experiment. In contrast, the ANI-2x potential has significantly higher errors, with MAE and RMSE errors twice as large as MACE-OFF23(M).
Most of the outliers predicted by MACE-OFF23(M) sit at the extreme ends of the spectrum. We found that the densities of low boiling point ether-containing compounds were under-predicted, likely because they begin to boil off at the simulation temperature, since the potential marginally under-predicts their boiling points. At the top end of the density range, the dibromo compounds were also under-predicted. This is perhaps unsurprising since there are only 84 configurations in the training set containing dimers of dibromo compounds.
@@
Interestingly, we found that MACE-OFF23(S) systematically overpredicts the experimental densities (SI Table 4) with a MAE of 0.23 g/cm3. This is likely the result of the slightly higher energy error and the shorter cutoff radius. In particular, the contribution of the intermolecular interactions to the overall energy is much smaller than the intramolecular interactions making the fractional error in them much larger. This is demonstrated in SI Table 3, where the test errors are decomposed into inter- and intramolecular contributions. We find that the medium and large models have significantly lower intermolecular force errors.
@@
We further investigate the performance of MACE-OFF23(M) on heats of vaporization. The predictions correlate strongly with the experimental data ($r^{2}$=0.88), however there is a systematic offset of approximately 2 kcal/mol. This may result from an under-prediction of the total intermolecular interactions originating from two sources: those that appear in molecular dynamics simulations outside the effective cutoff of the MACE model (c.f. classical simulations, which typically employ long-range corrections to the total energy [80]), and those that the model has not learned from the limited set of intermolecular interactions in the training set.
@@
Interestingly, for MACE-OFF23(S), this offset disappears (see SI Table 5), however this is likely due to the unphysically high predicted densities, resulting in a cancellation of errors through overprediction of the magnitude of intermolecular interactions. Nevertheless, these data indicate that, given a reasonable receptive field and a sufficiently expressive model, it is possible to fit a local force field that captures the interactions required to recover experimental properties.
III.6 Peptide Simulations

### v3

@@ III.1 Extended SPICE Test Set
First, we look at the pointwise errors of the energy and force predictions on a held-out test set for each of the three baseline MACE-OFF23 models. Figure 1 shows the per-atom energy, force and intermolecular force root-mean-square errors (full statistics are provided in SI Table 4). As the size of the model increases, the models gradually become more accurate, with the large model generally achieving errors of around 0.5–1.0 meV/atom and 15–20 meV/Å, well below the 1 kcal/mol (43 meV) chemical accuracy limit for the drug-like organic molecules studied here.
Figure 1: Test set root mean square errors (RMSE). Errors in the MACE-OFF23 models compared to the underlying DFT reference data, highlighting the relative accuracy of the three models. Bottom panels show specifically inter-molecular force errors compared to overall DFT inter-molecular force magnitudes (RMS) and note the absence of results for subsets that are composed of single molecules (i.e. DES370K Monomers, Dipeptides, QMugs, Tripeptides). Note that for subsets comprising only single molecule configurations (i.e. DES370K Monomers, Dipeptides, QMugs, Tripeptides) inter-molecular contributions are expected to be zero. The slight deviation from zero arises because DFT forces do not obey translational and rotational symmetries with sufficient accuracy, while MACE models do (check SI section B for details).
The last column represents an extrapolation task, the training set contains only dipeptides, and this test set looks at tripeptides, indicating that the models are able to extrapolate to larger fragments with more complex interactions, with no loss of accuracy.
The bottom panel of Figure 1 shows specifically the intermolecular force errors, which were obtained by separating the force contributions to molecular translations and rotations (see SI and Ref. [65] for details). These interactions are crucial as they underpin the thermodynamics and transport properties in the organic condensed phase. MACE-OFF23(M) and MACE-OFF23(L) both yield similar errors of around 5–15 meV/Å, which are about 1.5–3 times smaller than total force errors, and about 5–10% of the typical DFT intermolecular force magnitude (RMS). For context, relative intramolecular force errors, which are roughly equivalent to total force errors, are around 1–2%. The latter are routinely used across the literature to benchmark ML models, however, we advocate that intermolecular RMSE errors are a more appropriate metric to assess model accuracy for organic condensed phase applications. The results here demonstrate the remarkable accuracy of MACE models in predicting these subtle but crucial interactions. We will investigate how these energy and force errors translate into condensed phase physical property predictions in later Sections.
Figure 2: Test set prediction and root mean square errors (RMSE). Violin plots of the reference (blue) and predicted (red) distributions of the dipole moment predicted using MACE-OFF23-$\mu$. The RMSEs are reported above the violin plots for the individual test sets.
We also assess the accuracy of the MACE-OFF23-$\mu$ model on the test sets, evaluating its effectiveness for predicting dipole moments and infrared spectra of organic molecules. Figure 2 summarizes this performance by examining the alignment between the violin distributions of the reference and predicted total dipole moment values, as well as the RMSEs. The small RMSEs and strong alignment of the distributions indicate high fidelity across the test datasets for which reference dipole moment values were available. These results suggest promising performance for zero-shot predictions of the dipole moments (and infrared spectra) of unseen organic molecules, as demonstrated in a recent benchmark on out-of-distribution drug-like molecules from the NIST database [66].
@@ III.2 Dihedral scans
Next, we evaluate the performance of the MACE-OFF23 models on dihedral scans of drug-like molecules. This task is routinely carried out using quantum mechanical methods to create reference data for parameterizing classical empirical force field dihedral terms [67]. The task is particularly challenging, as constrained geometry optimizations can be difficult to converge if the potential energy surface is not smooth, as has been observed for example for the ANI-1x potential [68].
@@ III.2.1 TorsionNet-500
The top panel of Figure 3 summarizes the results for the TorsionNet-500 dataset [70]. This dataset contains torsion drives of 500 different molecules, selected to cover a wide range of pharmaceutically relevant chemical space. The original data set was reported at the B3LYP/6-31G(d) level of DFT theory. For consistency with the SPICE dataset, we recreated the torsion profiles using the DFT setting of SPICE, which uses a higher level of theory and basis set (SI Section C).
The first panel shows an example of a torsion drive, indicating the complex energy profile that the MACE models are able to capture closely, including geometries far from equilibrium. The center panel shows the mean barrier height error of a number of representative models, comparing the Sage classical empirical force field [21], a semi-empirical quantum mechanical method GFN2-xTB [23], a recent transferable machine learning force field AIMNet2 [8], and the three MACE-OFF23 models. Again, systematic improvements in accuracy with MACE model size are observed, with medium and large models, in particular, achieving errors of around 0.25 kcal/mol compared to the reference method. The AIMNet2 model achieves comparable accuracy to the small MACE-OFF23 model. A similar conclusion can be drawn from the comparison of the molecular geometries by looking at the root mean squared deviation of the atomic positions averaged over the full torsion scans, as indicated by the top right panel of Figure 3. It shows that MACE-optimized geometries have a deviation of about 0.025 Å, meaning they are almost indistinguishable from DFT-optimized structures. It is important to note that the other models were trained to different levels of DFT, which might also contribute to the observed differences.
@@ III.2.2 Biaryl fragments
The DFT torsion drives achieve a mean barrier height error of 0.2 kcal/mol with respect to coupled cluster data, and is therefore the best result theoretically possible for MACE-OFF23 models which were trained on the same DFT level of theory. We find that the large model comes close to this, with a mean absolute error of 0.3 kcal/mol. The medium and small MACE models have barrier height errors of 0.4 and 0.5 kcal/mol, respectively. Remarkably, the small MACE model is significantly more accurate than the next-best non-DFT methods, ANI-1ccx and GFN2-xTB, as illustrated in the bottom-center plot of Figure 3. In particular, unlike MACE-OFF23, coupled cluster reference calculations were used to parameterize the ANI-1ccx and GFN2 models. In the bottom right, we also show the cumulative error distributions to verify that the MACE-OFF23 barrier height errors are not only accurate on average but are also robust, with essentially no outliers.
III.2.3 Infrared spectroscopy of a drug molecule
To assess the accuracy of the MACE-OFF23 and MACE-OFF23-$\mu$ models for finite temperature properties of molecules, we study the infrared spectrum of an isolated paracetamol molecule at room temperature. This is a stringent test, as reproducing the line positions and their relative intensities requires an accurate treatment of both intramolecular interactions and the dipole moment in out-of-equilibrium configurations. As shown in Figure 4, we estimate the classical and approximate quantum IR spectra [71] using the MACE-OFF23(M) and MACE-OFF23-$\mu$ models, and compare them directly with experimental data at 293 K [72]. For low and intermediate frequencies (up to around 2000 cm-1), the classical and quantum spectra agree with each other and with the experimental data, allowing identification of all collective modes. For the high-frequency O–H and N–H modes, we observe that the classical spectra are shifted to the blue due to the absence of anharmonic zero-point fluctuations, which typically soften these bonds. A quantum treatment of nuclear dynamics results in good agreement with the experimental data, with the exception of a net blue shift, likely due to the limitations of the reference density functional. We also note a slight discrepancy in the relative intensities of the high-frequency stretching with respect to the lower-frequency modes. This could result from a limitation of the reference density functional or the MACE-OFF23-$\mu$ model, which we aim to investigate in a future study.
Figure 4: Infrared (IR) spectrum of paracetamol molecule. The classical (blue dashed) and quantum (blue solid) IR spectrum at 293 K using the MACE-OFF23(M) and MACE-OFF23-$\mu$ models by propagating the system and estimating the time correlation function of the time derivative of the total dipole moment. Quantum nuclear effects are incorporated using path-integral coarse-grained simulations (PIGS) [71]. The experimental data are taken from Ref. 72.
@@ III.3 Molecular crystals
In the following, we demonstrate the ability of the MACE-OFF force fields to simulate the vibrational and thermal properties of molecular crystals.
@@ III.3.1 Vibrational spectroscopy of paracetamol
Here, we test the ability of the MACE-OFF23(S) model to predict the Raman spectrum of the “Form II” polymorph of paracetamol. To compare MACE-OFF23(S) directly with experiments, we rigorously incorporate quantum nuclear and non-Condon effects using our ML-aided framework [73]. We incorporate quantum nuclear effects by fitting an effective potential energy surface [71] (using MACE) to calculate quantum nuclear corrections to the MACE-OFF23(S) model within the path-integral coarse-grained simulations (PIGS) method. To incorporate non-Condon effects, we fit a separate equivariant MACE model to the first-principles polarizability of paracetamol polymorphs (data taken from Refs. 74, 75). The remaining steps of our quantum nuclear simulations mimic an entirely classical calculation involving a $NVE$ molecular dynamics simulation, prediction of the isotropic and anisotropic components of the polarizability tensor, and calculation of their time correlation functions. For further details we refer the reader to Ref. 73.
@@ III.3.2 Lattice enthalphies
Figure 6 compares the predicted sublimation enthalpies with the experimentally measured ones. This task is often used to test various DFT functionals. Since the ωB97M-D3(BJ) functional used to parameterize MACE-OFF23 does not have a periodic implementation, an estimate of the highest achievable accuracy is not available. The figure shows that the three MACE models are able to capture trends and improve significantly compared to the ANI-2x model. The large MACE-OFF23 model achieves a mean absolute error of just 1.7 kcal/mol, which is comparable to the errors of several different dispersion-corrected density functionals for a fraction of the computational cost [78].
@@ III.4 Water structure and dynamics
A key requirement for bio-organic force fields is an accurate description of water. Several empirical water models have been developed for this purpose, the most common for all-atom biomolecular simulations being TIP3P. Figure 7 shows that MACE-OFF23(M) correctly predicts the radial distribution function (RDF) of water, performing comparably with both TIP3P and the highly accurate MB-pol water model, whose RDF is indistinguishable from experiment [80]. MACE-OFF23(S) gives a qualitatively similar RDF, with one noteable shift in the position of the second peak in the $g_{OO}(r)$ peak from 4.5 Å to 4 Å. This performance is consistent with the results discussed in more detail in section III.6. These interactions are modelled in a purely local way by MACE, having been trained on small water clusters containing up to 50 water molecules. This result further emphasizes the ability of the model to generalize to bulk simulations, even though the training data only contained non-periodic clusters. In contrast, ANI-2X shows a significant over-structuring of the RDF, likely due to the lack of dispersion correction in the underlying quantum mechanical method used to label the training set.
We further calculated the temperature dependence of the density using NPT molecular dynamics simulations, as shown in Figure 9. This is a highly sensitive test for the accuracy of intermolecular interactions [65]. Both the small and medium models over-predict the density to a small extent, which likely originates from the underlying DFT functional or the lack of periodic training data.
@@
Figure 7: Water structure and dynamics. Top: IR, isotropic Raman and anisotropic Raman spectra of water at ambient temperature and density using the MACE-OFF23(S) model for the potential energy surface and a single MACE model of water’s dipole and polarizability [73]. The black curve represents the experiment [82]. The dashed blue curve is obtained from standard (classical) MD, while the solid curve incorporates quantum effects using the PIGS method [71]. Bottom: Radial distribution function of MACE, ANI and TIP3P models. MB-pol traces are reproduced from Ref. [80].
III.5 Condensed phase properties of molecular liquids
Small molecule empirical force fields are fitted to reproduce experimental condensed phase properties including densities and heats of vaporization. For machine learning force fields, this task represents a greater challenge, since the potential is typically trained only using QM reference data for isolated molecules and molecular dimers. From these ab initio data, the models learn the long-range interactions necessary to reproduce bulk properties without ever fitting to them.
@@ III.5 Condensed phase properties of molecular liquids
To investigate the performance of the MACE-OFF23 models in the condensed phase, we selected a benchmark set of 109 molecules from Ref. [83], containing a representative set of functional groups relevant to medicinal chemistry and biology. MACE-OFF23(M) has a larger receptive field and is therefore better suited for reproducing experimental observables most affected by intermolecular interactions. In general, we found that the small model gave a reasonable description in the NVT ensemble, while the medium model is also able to accurately model liquids in NPT simulations, as detailed below. For the liquid boxes containing approximately 600 atoms, MACE-OFF23(S) achieved a throughput of $2.1\times 10^{6}$ steps/day via OpenMM on a single 80GB NVIDIA A100 GPU, whilst the medium model achieved a throughput of $1.1\times 10^{6}$ steps/day. Meanwhile MACE-OFF23(L) achieved a throughput of $2.8\times 10^{5}$ steps/day on the same system, highlighting the significant overhead compared to the other two models. See section III.7 for further discussion of the computational performance.
III.5.1 Organic liquids
@@
First, we investigate the predictions of molecular liquid densities under ambient conditions using the medium MACE model shown on the top panel of Figure 8. MACE-OFF23(M) achieves a MAE of 0.09 g/cm3, indicating that it can generalize from the intermolecular interactions between molecular dimers in the training set to the condensed phase, retaining a good correlation with experiment. In contrast, the ANI-2x potential has significantly higher errors, with MAE and RMSE errors twice as large as MACE-OFF23(M).
The vast majority of the outliers in the density predictions can be assigned as either low boiling-point ether-containing compounds, polychlorinated hydrocarbons or dibromo-containing compounds. In the first case, this may be due to a marginal under-prediction of the boiling point, since these compounds were simulated at 10 K below their tabulated boiling points. In the latter two cases, this functional group dependence suggests these nonbonded interactions are insufficiently represented in the present training set, which contains, for example, only 84 examples of dibromo-containing dimers. This highlights the need for additional coverage of specific functional groups within SPICE (for example, the recent AIMNet2 dataset includes 20 million configurations to cover 14 chemical elements [8]).
@@
Interestingly, we found that MACE-OFF23(S) systematically overpredicts the experimental densities (SI Table 6) with a MAE of 0.23 g/cm3. This is likely the result of the slightly higher energy error and the shorter cutoff radius. In particular, the contribution of the intermolecular interactions to the overall energy is much smaller than the intramolecular interactions making the fractional error in them much larger. This is demonstrated in SI Table 4, where the test errors are decomposed into inter- and intramolecular contributions. We find that the medium and large models have significantly lower intermolecular force errors. Whilst we see MACE-OFF23(L) as a replacement for semi-empirical QM due to its relative expense compared to the small and medium models, we nonetheless benchmarked the large model on a small selection of condensed phase systems. We saw comparable performance to the medium model, and in particular observed that compounds that had a high prediction error under the medium model had a similar error under the large model. This indicates that this error is driven primarily by the dataset rather than by the size of the model, as previously hypothesized.
@@
We further investigate the performance of MACE-OFF23(M) on heats of vaporization. The predictions correlate strongly with the experimental data); however, there is a systematic offset of approximately 2 kcal/mol. We hypothesized that this may result from an underprediction of the total intermolecular interactions originating from two sources: those that appear in molecular dynamics simulations outside the effective cutoff of the MACE model (compare classical simulations, which typically employ long-range corrections to the total energy [84]), and those that the model has not learned from the limited set of intermolecular interactions in the training set.
To investigate this further, we reran the calculations with MACE-OFF24(M) (see SI), which employs an extended cutoff of 6 Å and has also been trained on additional nonbonded data. We observe a modest decrease in MAE from 2.18 to 1.75 kcal/mol, as well as a small decrease in the accuracy of liquid density predictions. This further reinforces the hypothesis that the poorly predicted densities are a result of under-representation of the functional groups, since the additional training sets do little to increase sampling of these problematic functional groups.
@@
Interestingly, for MACE-OFF23(S), this offset disappears (see SI Table 7), however this is likely due to the overpredicted densities, resulting in a cancellation of errors through overprediction of the magnitude of intermolecular interactions.
Nevertheless, these data indicate that, given a reasonable receptive field and a sufficiently expressive model, it is possible to fit a local force field that captures the interactions required to recover experimental properties.
III.6 Biomolecular simulations